OpenStudy (anonymous):

Find x 4(5x+6)+2x=9x+12+13x

OpenStudy (mertsj):

20x+24+2x=22x+12 22x+24=22x+12 24=12 Oh no!!! That is not true. There is no solution.

OpenStudy (anonymous):

i knew it!! Thank you very much!!

OpenStudy (karatechopper):

merts u r so smart!

OpenStudy (mertsj):

Not as smart as satellite

OpenStudy (anonymous):

lol

OpenStudy (karatechopper):

yea but he is like college math or something higher! what grade r u in merts?

OpenStudy (mertsj):

I am not in school.

OpenStudy (anonymous):

to let you all know, im in middle school..

OpenStudy (anonymous):

bye thx

OpenStudy (karatechopper):

i am in middle school too. merts so u r an adult with a job?

OpenStudy (mertsj):

I do not have a job.

OpenStudy (mertsj):

I live in Obamaville

OpenStudy (karatechopper):

?